OpenStudy (anonymous):

Use a paragraph proof to prove the alternate interior angles theorem. Given: Prove: Alternate interior angles are congruent

OpenStudy (anonymous):

Given: AB ll CD

OpenStudy (anonymous):

@ParthKohli Any clue how to do this? haha><

Parth (parthkohli):

Hmm I do know how.

Parth (parthkohli):

It's indirect relationship.

OpenStudy (anonymous):

If you'd rather help me with this one: Given: 7x + 2 = 2 + 5x – 6 Prove: x = –3 Given the equation 7x + 2 = 2 + 5x – 6, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 7x + 2 = 2 – 6 + 5x. Then use the ____________________ to group the like terms. This gives the equation 7x + 2 = (2 – 6) + 5x. Next, combine like terms to get the equation 7x + 2 = –4 + 5x. Use the subtraction property of equality to subtract 5x from both sides of the equation. This gives the equation 2x + 2 = –4. Then use the subtraction property of equality again to subtract 2 from both sides of the equation. This gives the equation 2x = –6. Finally, use the division property of equality to divide both sides of the equation by 2 to give a final solution of x = –3. Therefore, given the equation 7x + 2 = 2 + 5x – 6, x is equal to –3. Which justification was left out of the paragraph proof above? By all means, please do! lol

OpenStudy (anonymous):

I'll only ask for help on one and try to figure the other one outt

Parth (parthkohli):

 The angle in black is vertical and the new angle formed is corresponding to the other angle.